Problem 18 Use the Divergence Theorem to calculate the surface integral || FdS , F(x,y,z) =< x²yz,xy-z, xyz? > S is the surfaShow Sketch and all steps.
Answer #1

Problem 18:- Given Vector E= <ayz, ny, ry zry.. . Given S is the surface of the boni enclosed by the planas . 1-0, 4-a, yao,10 b - on the face : OFF We have ñ= -2, x=0, ds=dydz. . . 9:07b, 3:0-e... .: S Funds = gb sp Cangasit (kg)]+(9 357) ).0)dy doAloed on the face sv. Antes We have 306, d9= drdy. ñ= 2:07a, 4:0-56. Noro Decade a S Sonygos dady. no yed so share y dady = cBy By Game Fido como Gauss Divergence theorem. I Founds = { divendu Casanoſt rutit (again) Sina F = Guya) + (499)ſt (1997) €